Part A Express your ansre usinwsonn fo Part B t the car contnss to decelerate at this rate, how much more time is required for it to stop? Express your anseer to two significent figures and Include the appropriate units. lfthe car co thues to decelerate ลthis rate, how far does it go? Find the total dstance
Answer #1

a)

Given

Vo = 86 Km/h =86*1000/3600 =23.89 m/s

V=59*1000/3600 =16.39 m/s

radius

r=d/2 =0.82/2 =0.41 m

angular speed

Wo=Vo/r =23.89/0.41=58.2656 rad/s

W=V/r =16.39/0.41 =39.973 rad/s

From

omega ^{2}=omega _{o}^{2}+2alpha Delta heta

39.973^{2}=58.2656^{2}+2alpha imes (95 imes 2pi )

alpha =-1.5053 rad/s^{2}

b)

From

omega =omega _{o}+alpha t

39.973 =58.2656+(-1.5053)t

t=12.15 s

c)

angular displacement before the car stops ,when the car stops final angular speed is zero i.e W=0

0 = 58.26562 + 2(-1.5053)▲θ

Delta heta =1127.64 rad

Total linear displacement

X=rDelta heta =0.41 imes 1127.64=462.33 m